### v1

# Solver-in-the-Loop: MDP-Based Benchmarks for Self-Correction and Behavioral Rationality in Operations Research

Abstract
[…]
Keywords: Operations Research, Large Language Models, Self-Correction, Reinforcement Learning, Benchmark
1 Introduction
1.1 From Translation to Debugging
When a linear program returns Infeasible, the real work begins. An analyst must examine the Irreducible Infeasible Subsystem (IIS)—the minimal subset of constraints that cannot be simultaneously satisfied—diagnose the root cause, and systematically repair the formulation. This iterative debugging loop is where OR expertise manifests.
Yet existing benchmarks evaluate LLMs on Operations Research as one-shot translation: given a problem description, generate solver code. This paradigm ignores the debugging process central to real OR practice. Unlike generic error messages, IIS provides a minimal certificate of infeasibility, enabling targeted, interpretable repairs—precisely the kind of structured feedback that should enable targeted self-correction.
Figure 1: Evaluation paradigms compared. Top: Static translation benchmarks evaluate one-shot code generation with no execution feedback. Bottom: Our solver-in-the-loop approach enables iterative self-correction through IIS feedback.
1.2 Self-Correction in Structured Domains
Recent work showed that 64.5% of LLM errors result when models fail to self-correct (Tie et al., 2025). However, CorrectBench focuses on general programming tasks and omits the OR domain, where structured self-correction is uniquely suited. First, solvers provide deterministic feedback—precise, verifiable signals such as IIS, slack values, and objective bounds. Second, verifiable ground truth enables mathematical checking of optimal solutions. Third, the interpretable process means the diagnostic reasoning chain admits structured evaluation.
This combination—deterministic oracle, verifiable outcomes, structured process—makes OR an ideal testbed for studying self-correction. The solver-in-the-loop paradigm forces agents to reason from IIS feedback, enabling systematic hypothesis refinement rather than blind trial-and-error.
1.3 Behavioral Rationality in Operations
While debugging addresses upstream formulation errors, operational decisions face a distinct downstream challenge. Concurrent work on AIM-Bench (Zhao et al., 2025) revealed systematic behavioral biases in LLM inventory managers: a “pull-to-center” tendency where models over-order when optimal quantity is low and under-order when it is high. This bias persists across model scales, raising concerns about deploying LLMs in high-stakes operations management.
1.4 Contributions
We introduce two benchmarks and a training framework:
  1. 1.
    Two complementary benchmarks: OR-Debug-Bench ($\scriptstyle\sim$5,000 problems, 9 error types) evaluates iterative debugging with solver-verified ground truth; OR-Bias-Bench (2,000 instances, ID/OOD splits) evaluates behavioral rationality against closed-form optimal policies.
  2. 2.
    RLVR training with process rewards: Domain-specific GRPO training with outcome verification, diagnostic accuracy, and faithfulness penalties enables an 8B model to surpass frontier APIs: +9.1% RR@5 (95.3% vs 86.2%) and +14.6% DA (62.4% vs 47.8%).
  3. 3.
[…]
  4. 4.
    Evaluation spanning 26 models: 12,000+ samples across frontier APIs and local variants, with public release of benchmarks, training pipelines, and evaluation code.
Table 1: Key results summary.
Finding Metric Result
8B surpasses frontier APIs RR@5 95.3% vs 86.2%
Diagnostic accuracy gain DA 62.4% vs 47.8%
Efficiency improvement Steps 2.25 vs 3.78
OOD generalization Bias drift -9.6% (only negative)
[…]
2 Benchmark Setup
Figure 2: Two-phase benchmark framework. Phase I (OR-Debug-Bench): Iterative debugging where the agent receives Gurobi IIS feedback and repairs infeasible code. Phase II (OR-Bias-Bench): Inventory decision-making verified against the closed-form optimal policy $Q^{*}=F^{-1}(\text{CR})$.
2.1 OR-Debug-Bench: Data Organization & Metrics
Table 2: Benchmark positioning. OR-Debug-Bench is the first to evaluate iterative debugging with deterministic oracle feedback.
Benchmark Year Task Oracle Multi-step Self-Corr.
[…]
OptiBench (Yang et al., 2024) 2024 Formulation
MAMO (Huang et al., 2024) 2024 Complex LP
ORLM (Huang et al., 2025) 2025 Formulation
AIM-Bench (Zhao et al., 2025) 2025 Inventory Closed-form
SWE-bench (Jimenez et al., 2024) 2024 Code Debug Unit Tests $\checkmark$ Limited
CorrectBench (Tie et al., 2025) 2025 Self-Corr. General
OR-Debug-Bench 2026 Debugging Gurobi IIS $\checkmark$ $\checkmark$
OR-Bias-Bench 2026 Decision Closed-form $\checkmark$
Each problem consists of: (1) a natural language description paired with sabotaged Gurobi code producing Infeasible, (2) ground truth including original code, IIS constraints, and target error type, and (3) Gurobi 11.0 IIS computation as the oracle. Figure 3 illustrates a complete episode where the agent diagnoses an IIS conflict and repairs the model in two steps.
Table 3: OR-Debug-Bench benchmark statistics and data splits.
Attribute Value
Dataset Structure
Total Problems $\scriptstyle\sim$5,000
Error Types 9 (A–I)
IIS Range 1–11 constraints
Difficulty Levels Easy / Medium / Hard / Expert
Data Splits
Training (SFT) 696 trajectories
Training (RL) 300 prompts
Evaluation (Stratified) 450 (50 per type)
Max Steps per Episode 50
Figure 3: Example OR-Debug-Bench episode. Left: The agent receives a sabotaged LP where minimum requirements ($60+50+0=110$) exceed capacity (100100). Right: The agent (1) attempts optimization, (2) computes IIS, (3) reasons about the conflict, (4) relaxes the key constraint, and (5) achieves Optimal status in 2 repair steps.
Training-Evaluation Split. Training uses <<20% of total problems (996 of $\scriptstyle\sim$5,000); evaluation samples are drawn from a held-out pool excluding all training IDs. For OR-Bias-Bench, training uses 900 samples across three curriculum stages with CR range [0.3, 0.7], while evaluation covers broader ranges: ID [0.05, 0.95] and OOD [0.10, 0.89].
Metrics. RR@k (Recovery Rate) measures percentage reaching Optimal within $k$ steps. DA (Diagnostic Accuracy) = $|\mathcal{D}\cap\text{IIS}{}|/|\text{IIS}{}|$ measures root cause identification. OP (Optimality Preservation) = $1-|\Delta\text{obj}|/|\text{obj}_{\text{orig}}|$ measures solution quality.
2.2 OR-Bias-Bench: Data Organization & Metrics
OR-Bias-Bench comprises 2,000 newsvendor instances (1,000 ID + 1,000 OOD) evaluating inventory decisions where ground truth $Q^{*}=F^{-1}(\text{CR})$ is a closed-form analytical solution—no solver interaction occurs.
Table 4: OR-Bias-Bench data splits used in experiments. The benchmark contains 2,000 instances (1,000 ID + 1,000 OOD); we evaluate on stratified subsets.
Attribute Train ID Eval OOD Eval
Samples 900 400 200
CR Range [0.3, 0.7] [0.05, 0.95] [0.10, 0.89]
Demand Dist. $\mathcal{N}(\mu,\sigma)$ $\mathcal{N}(\mu,\sigma)$ $\mathcal{N}(\mu,\sigma)$
$\mu$ Range [50, 200] [50, 200] [50, 200]
$\sigma$ Range [10, 50] [10, 50] [10, 50]
[…]
Table 5: Comparison of the two benchmarks: OR-Debug-Bench targets upstream model repair through iterative solver interaction, while OR-Bias-Bench evaluates downstream decision-making against analytical ground truth.
Aspect OR-Debug-Bench OR-Bias-Bench
Domain Mathematical Programming Operations Management
Task Debug infeasible code Inventory decision
Oracle Gurobi IIS feedback Closed-form $Q^{*}=F^{-1}(\text{CR})$
Interaction Multi-step (up to 20) Single-shot
LLM Challenge Error parsing, code repair Cognitive bias mitigation
Key Metrics RR@k, DA, OP Rationality, Bias Diff
2.3 Evaluation Protocol
For OR-Debug-Bench, each episode proceeds: (1) agent receives initial state (problem, sabotaged code, Infeasible, IIS), (2) agent outputs action, (3) environment executes and returns new state, (4) repeat until Optimal or max steps. For OR-Bias-Bench, the agent outputs a single order quantity $Q$.
We evaluate 26 models on 450 stratified OR-Debug-Bench samples per model and 600 OR-Bias-Bench samples (400 ID + 200 OOD) spanning frontier APIs and local variants.
3 Benchmark Construction
3.1 Saboteur-based Problem Generation
We design a “saboteur” pipeline that injects controlled errors into valid linear programs while maintaining verifiable ground truth. Each generated error must: (1) produce a verifiable Infeasible status, (2) yield a non-empty IIS containing the sabotaged constraint, and (3) have a unique ground-truth fix restoring Optimal status.
Table 6: Error type taxonomy for OR-Debug-Bench.
Type Name IIS Size Difficulty
A Direction Flip 2–3 Easy
B RHS Miscalculation 3–5 Medium
C Upper Bound Conflict 2–3 Easy
D Lower Bound Conflict 2–4 Easy
E Resource Over-allocation 5–8 Hard
F Capacity Violation 5–7 Hard
G Flow Imbalance 6–10 Hard
H Multi-constraint Conflict 8–12 Expert
I Composite Error 10–15 Expert
Generation Pipeline. The pipeline operates in four stages: (1) source selection from a feasible LP pool, (2) sabotage application using type-specific corruption, (3) verification via Gurobi IIS computation, and (4) oracle labeling for evaluation. Of the initial candidate pool, 87% pass all validation checks on first generation; the remainder require at most two iterations. Full algorithmic details appear in Appendix C.1.
Robust Injection. Reliable error injection uses adaptive methods: Type A (direction flip) uses slack-based constraint selection to improve success from 30% to 95%; Type C (upper bound conflict) uses a 4-tier fallback strategy achieving 72% success. Complete algorithms appear in Appendix C.8.
Table 7: Difficulty calibration for OR-Debug-Bench. Each level targets specific IIS sizes and expected SFT baseline performance.
Level IIS Range Error Types Target SFT RR@5
Easy 2–4 A, B, C, D $\geq$90%
Hard 5–10 E, F, G 70–85%
Expert 8–15 H, I 45–70%
3.2 Anti-Pattern Measures
Three mechanisms prevent pattern-matching shortcuts. First, randomized naming in Types G–I uses UUID-based identifiers to prevent models from exploiting semantic name patterns. Second, hidden dependencies (Type F) create scenarios where the IIS reveals a symptom constraint while the root cause lies elsewhere. Third, cascading conflicts (Type G) require multi-step reasoning: fixing the primary conflict reveals a secondary one. Additional details appear in Appendix C.10.
3.3 Newsvendor Problem Generation
For OR-Bias-Bench, we generate newsvendor scenarios with controlled critical ratios:
$Q^{*}=\mu+\sigma\cdot\Phi^{-1}(\text{CR}),\quad\text{CR}=\frac{p-c}{p-s}$ (1)
where $\Phi^{-1}$ is the standard normal inverse CDF, $\mu$ is mean demand, and $\sigma$ is standard deviation.
Stratified Sampling. We stratify by CR buckets: ID covers [0.05, 0.95] with 100+ samples per bucket; OOD covers [0.10, 0.89] testing intermediate-to-extreme values. The 4-level curriculum is detailed in Appendix C.3.
Table 8: Curriculum levels for OR-Bias-Bench. Each level targets specific bias phenomena with controlled CR ranges and prompt complexity.
Level CR Range Prompt Style Target Phenomenon
L1 $[0.4,0.6]$ Clean Foundation (neutral)
L2 $0.05,0.2)\cup(0.8,0.95]$ Clean Bias trigger (extreme)
L3 $[0.3,0.7]$ + Distractors Robustness test
L4 $[0.1,0.9]$ + Censored Expert inference
3.4 MDP Formulation
We formalize both benchmarks as Markov Decision Processes.
OR-Debug-Bench MDP. The state, action, and reward are:
$\mathcal{S}$ $=(\text{problem}_{\text{NL}},\text{code},\text{status},\text{iis},\text{history},t)$
$\mathcal{A}$ $=\{\textsc{Get\_IIS},\textsc{Check\_Slack},\textsc{Relax},$
$\quad\ \,\textsc{Drop},\textsc{Rewrite},\textsc{Submit}\}$ (2)
$\mathcal{R}$ $=0.5\cdot R_{\text{outcome}}+0.3\cdot R_{\text{diagnosis}}+0.2\cdot R_{\text{efficiency}}$
Actions are organized hierarchically: Diagnostic (information gathering), Repair (code modification), and Meta (episode control). The oracle is Gurobi IIS computation providing deterministic state transitions. Full specifications appear in Appendix D.1.
OR-Bias-Bench MDP. The bias evaluation uses a single-step MDP where $\mathcal{A}=\{Q\in\real^{+}:Q\geq 0\}$ and the oracle is the closed-form solution in Eq. (1).
Diagnostic Accuracy (DA). We measure alignment between diagnosed constraints and ground truth:
$\text{DA}=\frac{|\text{diagnosed}\cap\text{IIS}_{\text{GT}}|}{|\text{IIS}_{\text{GT}}|}$ (3)
High RR@5 with low DA indicates “lucky” solutions that fix problems without understanding root causes.
4 Training Methods
Figure 4: Training pipeline overview. Track 1 trains the OR-Debug-Bench model: SFT on teacher trajectories followed by GRPO with composite reward and optional PRM supervision. Track 2 trains the OR-Bias-Bench model: SFT on rational responses followed by a three-stage curriculum (Extreme $\rightarrow$ Boundary $\rightarrow$ Full) that targets pull-to-center bias.
Figure 4 illustrates our two-track training approach. Both tracks start from Qwen3-8B and apply SFT, but diverge in their RL phases: OR-Debug-Bench uses GRPO with solver-verified rewards, while OR-Bias-Bench uses curriculum learning to overcome systematic bias. The following subsections detail each component.
4.1 Foundation Model Selection
We selected Qwen3-8B-Instruct based on a pilot study (100 samples) showing +41.9% post-SFT improvement headroom.
Table 9: Pilot study: foundation model screening on OR-Debug-Bench validation.
Model Base RR@5 +SFT RR@5 $\Delta$
Qwen3-8B 51.2% 93.1% +41.9%
Qwen3-8B achieved 93.1% RR@5 after SFT with efficient token usage (2,100 tokens/episode). See Appendix I for methodology.
4.2 SFT Data Collection
We collect successful debugging trajectories from three teacher models—GPT-5.2-chat (40%), o4-mini (35%), and DeepSeek-R1 (25%)—chosen for diverse reasoning strategies. Filtering retains high-quality trajectories:
$\text{success}=\text{True}\land\text{steps}\leq 5\land\text{DA}\geq 0.5$ (4)
This yields 696 of 1,247 trajectories (55.8% acceptance) averaging 2.3 steps with 68% diagnostic accuracy. For OR-Bias-Bench, we collect 500 rational responses. Training examples are shown in Appendix D.10.
4.3 GRPO Training with Composite Reward
Following DeepSeek-R1 (DeepSeek-AI, 2025), we use Group Relative Policy Optimization with KL removal ($\beta=0$), asymmetric clipping $[0.2,0.28]$, and LoRA ($r=16$, $\alpha=32$).
Composite Reward Function. The reward balances outcome verification, diagnostic quality, and efficiency:
$R=0.5\cdot R_{\text{outcome}}+0.3\cdot R_{\text{diagnosis}}+0.2\cdot R_{\text{efficiency}}$ (5)
where $R_{\text{outcome}}=+100$ for Optimal and 50-50 otherwise, $R_{\text{diagnosis}}=\text{DA}\cdot 100$, and $R_{\text{efficiency}}=-1$ per step. The 30% diagnostic weight addresses trial-and-error solutions achieving Optimal without correct reasoning. A faithfulness penalty (20-20) discourages repairs targeting non-IIS constraints—without this penalty, models often achieve Optimal through indirect fixes that mask the root cause. Training converges after 4 epochs with RR@5=95.0%. Full training curves appear in Appendix D.5.
4.4 Process Reward Model (PRM)
Outcome-based rewards provide sparse feedback. We train a PRM (Qwen3-8B with LoRA) to provide step-level supervision:
[…]
4.5 Curriculum Learning for Bias Mitigation
For OR-Bias-Bench, we use a three-stage curriculum targeting the pull-to-center bias:
Table 10: Three-stage curriculum for OR-Bias-Bench.
Stage Focus Samples CR Distribution
1 Direction Learning 200 Extreme (0.1, 0.9)
2 Boundary Refinement 300 Near-boundary
3 Full Distribution 400 [0.2, 0.8]
Stage 1 (Direction Learning): Extreme CR values (0.1, 0.9) teach directional sensitivity. Stage 2 (Boundary Refinement): Near-boundary values ([0.15, 0.25] and [0.75, 0.85]) refine magnitude estimation. Stage 3 (Full Distribution): Coverage across [0.2, 0.8] consolidates learning.
[…]
5 Experiments
5.1 Experimental Setup
[…]
5.2 Main Results
Table 11: OR-Debug-Bench results on representative models (450 stratified samples). Full results for all 26 models in Appendix E.
Model $R$ RR@5 DA Steps
Qwen3-8B-GRPO 100% 95.3% 62.4% 2.25
Qwen3-8B-Curriculum 100% 94.0% 61.7% 2.22
Qwen3-8B-DAPO 100% 93.8% 60.4% 2.31
Qwen3-8B-SFT 99.8% 93.1% 60.8% 2.34
o4-mini 97.8% 86.2% 47.8% 3.15
claude-sonnet-4 100% 86.2% 50.1% 3.71
o1 99.8% 82.9% 47.8% 3.78
gpt-5.2-chat 99.8% 81.8% 40.9% 3.72
gemini-2.5-flash 84.2% 70.7% 19.2% 3.23
Llama-3.3-70B 93.8% 60.9% 46.9% 4.81
DeepSeek-V3.2 99.3% 58.9% 44.8% 4.86
DeepSeek-R1 99.1% 56.7% 34.5% 5.08
[…]
The trained models use a “diagnose once, repair correctly” pattern: 1.3 diagnostic actions per episode vs 2.1 for API models, then targeted repairs. This reflects fundamentally different reasoning—systematic elimination vs trial-and-error.
Per-Error-Type Performance. Domain-specific training provides larger gains on harder problems: +13.3% on Types E–G (94.7% vs 81.4%) and +20.8% on Types H–I (89.1% vs 68.3%). All models achieve >>90% on Types A–D, confirming the baseline difficulty calibration. Full breakdown appears in Appendix E.2.
[…]
5.3 Bias Evaluation Results
Table 12: OR-Bias-Bench results (400 ID / 200 OOD samples). Bias = difference from rational ordering.
Model ID Bias OOD Bias $\Delta$ Status
claude-haiku-4.5 0.0% 3.6% +3.6% Best ID
Qwen3-8B-OM-SFT 4.9% 11.5% +6.6% OK
o4-mini 6.7% 7.7% +1.0% OK
Qwen3-8B-Curriculum 20.0% 10.4% -9.6% Best OOD
gpt-4.1 11.8% 15.4% +3.6% OK
Llama-3.3-70B 19.2% 12.5% -6.7% OK
gpt-5-mini 1.2% 53.3% +52.1% Degraded
[…]
5.4 Inference Scaling Analysis
Figure 5: Recovery Rate vs. attempt budget $k$. Qwen3-8B-GRPO achieves at $k=5$ (95.3%) what the best API model requires $k>10$ to approach.
[…]
5.5 Ablation Study
Table 13: Ablation study on OR-Debug-Bench (200-sample validation set).
[…]
SFT Baseline 91.5% 68.0%
+ GRPO 92.0% 66.0% +0.5%
+ Curriculum 95.0% 68.0% +3.5%
+ PRM 92.0% 72.7% +0.5%
Curriculum + GRPO (best) 95.3% 62.4% +3.8%
Key Ablation Findings: (1) Curriculum pre-training provides +3.5% RR@5, the largest single improvement; (2) PRM improves DA by +4.7% but trades off against RR@5 (-3%); (3) Curriculum synergizes with GRPO (+3.8% over SFT), providing favorable initialization for hard problems.
6 Related Work
OR Benchmarks. Table 2 compares OR-Debug-Bench and OR-Bias-Bench with existing OR benchmarks. NL4Opt (Ramamonjison et al., 2023) introduced the task of translating natural language to linear program formulations. Subsequent work explored both prompt-based approaches using in-context learning (Xiao et al., 2024; AhmadiTeshnizi et al., 2024; Bertsimas and Margaritis, 2024) and learning-based methods fine-tuning on domain data (Yang et al., 2024; Huang et al., 2025). Other recent efforts include ORQA (Mostajabdaveh et al., 2025) for OR reasoning evaluation, LLMOPT (Jiang et al., 2024), and agent-based systems (Zhang and Luo, 2025; Thind et al., 2025). OptMATH (Lu et al., 2025) proposed bidirectional data synthesis for scalable training data generation. MAMO (Huang et al., 2024) further contributed benchmarks spanning easy and complex LP instances with solver integration. PILOT-Bench (Ao et al., 2026b) evaluates LLM workflow execution under probabilistic tool failures and varying instruction quality. However, all existing benchmarks evaluate static, one-shot formulation accuracy: a model produces code, and evaluation checks correctness without iterative refinement or solver feedback. Recent work has extended formulation benchmarks to dynamic programming: DP-Bench (Zhou et al., 2025) introduces 132 textbook-level DP problems and shows that even SOTA models struggle with stochastic transitions, achieving only 59.1% accuracy. Our work addresses a complementary gap: evaluating how models respond to solver feedback and iteratively correct their formulations, focusing on LP infeasibility debugging rather than one-shot formulation.
Self-Correction in LLMs. While Chain-of-Thought prompting (Wei et al., 2022) improved LLM reasoning through step-by-step decomposition, it does not address error correction when intermediate steps fail. CorrectBench (Tie et al., 2025) provided the first systematic study of LLM self-correction, documenting that 64.5% of errors stem from correction failures rather than initial mistakes. However, CorrectBench focuses on general programming tasks and explicitly excludes the OR domain, where feedback is deterministic and mathematically verifiable. SWE-bench (Jimenez et al., 2024) evaluates code debugging through unit test feedback, but software testing fundamentally differs from mathematical optimization verification: unit tests sample behavior while solvers provide complete, deterministic feedback. Self-Refine (Madaan et al., 2023) demonstrated iterative refinement with self-generated feedback, and STaR (Zelikman et al., 2022) showed that models can bootstrap reasoning by learning from their own correct solutions. Kumar et al. (2024) further demonstrated that RL can explicitly train models for iterative self-correction. However, these approaches rely on heuristic quality signals rather than formal verification. Our benchmarks leverage Gurobi’s IIS computation as a noise-free oracle, enabling precise evaluation of diagnostic reasoning. For multi-agent debugging scenarios, AgentGit (Li et al., 2025) provides version control abstractions that could complement our single-agent evaluation framework.
RLVR and Process Supervision. Reinforcement Learning with Verifiable Rewards (RLVR) has emerged as a powerful paradigm for training reasoning models. Building on Proximal Policy Optimization (PPO) (Schulman et al., 2017) and instruction tuning with human feedback (Ouyang et al., 2022), alternative approaches include Direct Preference Optimization (DPO) (Rafailov et al., 2023) for offline alignment and Group Relative Policy Optimization (GRPO(Shao et al., 2024) for online training. DeepSeek-R1 (DeepSeek-AI, 2025) demonstrated that GRPO with outcome verification can induce sophisticated reasoning without explicit chain-of-thought supervision. DAPO (Yu et al., 2025) further improved RLVR systems with KL-penalty removal and asymmetric clipping. Tülu 3 (Lambert et al., 2025) extended this to a broader post-training pipeline. Recent analysis (Chu et al., 2025) reveals that while SFT tends to memorize training patterns, RL promotes generalization—a finding that motivates our two-stage training approach. For process supervision, “Let’s Verify Step by Step” (Lightman et al., 2024) introduced human-annotated step-level rewards, later automated by Math-Shepherd (Wang et al., 2024) through Monte Carlo estimation. BiPRM (Zhang et al., 2025a) proposed bidirectional verification for mathematical reasoning. PAVs (Setlur et al., 2025) formalized Process Advantage Verifiers for efficient alignment. Our PRM training builds on these foundations, adapting process supervision to the OR debugging domain where step-level progress can be automatically verified through IIS size reduction. Concurrent work OR-R1 (Ding et al., 2025) also explores test-time RL for OR modeling, though focusing on formulation rather than iterative debugging.
Verifiable Feedback in OR. Prior self-correction work relies on either self-generated feedback, which can be unreliable, or heuristic metrics like test pass rates, which sample rather than verify. Solver feedback differs fundamentally: IIS computation provides complete and deterministic information about the infeasibility structure. We adapt RLVR to the OR domain where: (1) the oracle provides verifiable rewards without human annotation, (2) step-level progress is measurable through IIS size reduction, and (3) diagnostic accuracy can be computed against ground-truth constraint labels—enabling fully automated training and evaluation pipelines.
[…]
Behavioral Rationality in LLMs. Recent work has examined LLMs as simulators of human decision-making in OM (Zhang et al., 2025b). AIM-Bench (Zhao et al., 2025) documented the “pull-to-center” phenomenon in LLM inventory decisions, where models regress toward mean predictions regardless of the optimal decision. This finding echoes classical behavioral OR on human decision-making under uncertainty (Schweitzer and Cachon, 2000). Our OR-Bias-Bench benchmark extends AIM-Bench with explicit ID/OOD splits and demonstrates that curriculum learning can mitigate these biases, achieving 48% bias reduction on OOD scenarios.
Connection to Human Behavioral Biases. The pull-to-center bias in LLMs mirrors documented human biases in newsvendor decisions (Schweitzer and Cachon, 2000; Kremer et al., 2010). Our curriculum training, which explicitly teaches directional sensitivity, can be viewed as a cognitive debiasing intervention for LLMs, analogous to decision support systems designed for human operators.
7 Conclusion
We introduced OR-Debug-Bench and OR-Bias-Bench, complementary benchmarks evaluating iterative self-correction and behavioral rationality rather than one-shot formulation. Through experiments on 12,000+ samples across 26 models, we demonstrate:
  1. 1.
[…]
  2. 2.
    Process evaluation reveals coincidentally correct repairs: DA distinguishes correct diagnostic reasoning from trial-and-error success.
  3. 3.
    Curriculum learning enables genuine generalization: Staged training is the only approach showing improved OOD performance (-9.6% bias drift).
  4. 4.
    PRM improves diagnostic quality: Step-level supervision increases DA by +4.7%.
Broader Significance. Although our tasks are structurally simple, the Action $\to$ Feedback $\to$ Plan Update loop is foundational for reliable agent deployment. Domain-specific training within deterministic verification loops demonstrates a scalable path toward reliable decision-making in mission-critical domains.
Future Work. Extensions include MIP/MINLP debugging, multi-period operations, RAG integration with OR knowledge bases, and multi-agent collaboration.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix Overview
This appendix provides supporting material organized as follows:
  • Appendix A: Execution examples comparing trained and baseline models
  • Appendix B: Task parameters, prompts, and evaluation criteria
  • Appendix C: Benchmark construction algorithms and validation
  • Appendix D: MDP formulation and training procedures
  • Appendix E: Complete evaluation results
  • Appendix H: Token efficiency and difficulty analysis
  • Appendix I: Base model selection and negative results
Appendix A Execution Examples
This appendix provides execution traces illustrating the difference between trained and baseline models on OR-Debug-Bench problems.
A.1 Success Case: Type C Error (Upper Bound Conflict)
Problem Description. A production planning LP where minimum requirements on individual products exceed the total capacity constraint:
m = gp.Model("production")
x = m.addVars(3, name="x")
m.addConstr(x[0] + x[1] + x[2] <= 100, name="c1_total")
m.addConstr(x[0] >= 60, name="c2_min_0")
m.addConstr(x[1] >= 50, name="c3_min_1")  # KEY
m.addConstr(x[2] >= 0, name="c4_min_2")
The minimum sum ($60+50+0=110$) exceeds capacity ($100$), causing infeasibility.
Initial State.
  • Status: Infeasible
  • IIS: {c1_total, c2_min_0, c3_min_1}
  • Ground Truth Fix: Relax c3_min_1 to $x[1]\geq 40$
Qwen-GRPO Trajectory (2 steps).
  1. 1.
    Step 1: Diagnostic Action
    • Action: Get_IIS
    • Response: “IIS contains c1_total, c2_min_0, c3_min_1. The minimum sum $(60+50+0=110)$ exceeds capacity $(100)$. Key constraint: c3_min_1 ($x[1]\geq 50$) is over-restrictive.”
    • Diagnosis: [c3_min_1] ✓ (matches ground truth)
  2. 2.
    Step 2: Repair Action
    • Action: Relax(c3_min_1, $x[1]\geq 40$)
    • Response: “Modified constraint allows sum =60+40+0=100==60+40+0=100= capacity. Model is now feasible.”
    • Status: Optimal
SFT Baseline Trajectory (5 steps).
  1. 1.
    Get_IIS $\rightarrow$ [c1, c2, c3]
  2. 2.
    Check_Slack(c1) $\rightarrow$ “c1 is binding”
  3. 3.
    Relax(c2) $\rightarrow$ Still Infeasible
  4. 4.
    Relax(c1) $\rightarrow$ Status changes but objective degraded
  5. 5.
    Rewrite(c3) $\rightarrow$ Finally Optimal
Analysis. The GRPO-trained model directly identifies the key constraint (c3_min_1) and applies a minimal fix in 2 steps. The SFT baseline explores multiple constraints through trial-and-error, taking 5 steps and initially relaxing the wrong constraint (c2).
A.2 Challenge Case: Composite Error (Type I)
Problem Description. A composite error with 12 constraints in the IIS, involving resource allocation, capacity, and flow balance conflicts simultaneously.
Model Behavior Comparison.
Table 14: Model behavior comparison on composite error case.
Model Steps DA Success Analysis
Qwen-GRPO 7 58% Yes Systematic decomposition
o4-mini 12 42% Yes Trial-and-error
gpt-5.2-chat 15 33% Yes Excessive exploration
gpt-4.1-mini 20 17% No Error cascade
Error Cascade Pattern (gpt-4.1-mini).
  • Steps 1–5: Correct diagnosis of constraints c3, c7
  • Step 6: Incorrect fix of c3 introduces new conflict
  • Steps 7–12: Attempts to fix cascading errors
  • Steps 13–18: Reverts and retries with different approach
  • Steps 19–20: Timeout without resolution
Analysis. API models often fail to reason about constraint interactions, leading to fixes that introduce new conflicts. The trained model learns to decompose complex IIS sets and address constraints in dependency order.
A.3 Available Actions
Action Description Modifies State
Get_IIS Compute Irreducible Infeasible Subsystem No
Check_Slack Get constraint slack values No
Check_Bound Get variable bound status No
[…]
Drop(c)(c) Remove constraint entirely Yes
[…]
Submit Submit current model for evaluation Yes
A.4 Additional Success Examples
Example: Type E (Resource Over-allocation). A transportation problem where demand exceeds supply:
# Problem setup
m = gp.Model("transport")
supply = m.addVars(3, name="s")
demand = m.addVars(4, name="d")
# Supply constraints (total = 100)
m.addConstr(supply[0] <= 40, "s0_cap")
m.addConstr(supply[1] <= 35, "s1_cap")
m.addConstr(supply[2] <= 25, "s2_cap")
# Demand constraints (total = 115, infeasible!)
m.addConstr(demand[0] >= 30, "d0_min")
m.addConstr(demand[1] >= 35, "d1_min")  # KEY
m.addConstr(demand[2] >= 25, "d2_min")
m.addConstr(demand[3] >= 25, "d3_min")
# Flow balance
m.addConstr(quicksum(supply) == quicksum(demand))
IIS: {s0_cap, s1_cap, s2_cap, d0_min, d1_min, flow_balance}
Qwen3-8B-GRPO Solution (2 steps):
  1. 1.
    Get_IIS: “Total supply (100) is less than total demand (115). The largest demand constraint (d1_min = 35) should be reduced.”
  2. 2.
    Relax(d1_min, -15): “Reducing demand 1 from 35 to 20 balances supply and demand.” $\rightarrow$ Optimal
gpt-4.1 Solution (6 steps):
  1. 1.
    Get_IIS $\rightarrow$ Lists all constraints
  2. 2.
    Relax(s0_cap, +10) $\rightarrow$ Still Infeasible
  3. 3.
    Relax(s1_cap, +10) $\rightarrow$ Still Infeasible
  4. 4.
    Check_Slack(all) $\rightarrow$ Sees all slacks are 0
  5. 5.
    Relax(d3_min, -10) $\rightarrow$ Still Infeasible
  6. 6.
    Relax(d1_min, -15) $\rightarrow$ Optimal
Analysis: The trained model directly identifies that demand constraints are easier to adjust than capacity constraints (a domain-specific heuristic) and targets the largest demand. The API model attempts supply increases first, which don’t resolve the fundamental imbalance.
A.5 OR-Bias-Bench Examples
ID Scenario: CR = 0.5 (Balanced).
Price: \$60, Cost: \$30, Salvage: \$0
Mean demand: 100, Std: 20
CR = (60-30)/(60-0) = 0.5
[…]
Model responses:
  • gpt-5-mini: $Q=100$ (Correct, bias = 0%)
  • Qwen3-8B-Curriculum: $Q=100$ (Correct)
OOD Scenario: CR = 0.1 (Low margin).
Price: \$55, Cost: \$50, Salvage: \$5
Mean demand: 100, Std: 20
CR = (55-50)/(55-5) = 0.1
[…]
Model responses:
  • gpt-5-mini: $Q=95$ (Over-order, bias = +28%)
  • Qwen3-8B-Curriculum: $Q=76$ (Correct, bias = +2%)
Analysis: At extreme CR values, gpt-5-mini exhibits severe pull-to-center bias, ordering near the mean despite the optimal quantity being 32% below. The curriculum-trained model correctly adjusts its order downward, demonstrating learned sensitivity to the critical ratio.
Appendix B Task Function Components
This appendix details the task parameters, prompts, and evaluation criteria used in OR-Debug-Bench.
B.1 Task and Tool Details
Table 15 lists the key parameters governing the OR-Debug-Bench environment.
Table 15: Task parameters for OR-Debug-Bench.
Parameter Value Description
max_steps 50 Maximum MDP steps before timeout
timeout 10s Per-step solver timeout
iis_method minimal Gurobi IIS computation method
Table 16 describes the available actions and their properties. Diagnostic actions gather information without modifying the model state, while repair actions apply changes.
Table 16: Action space for OR-Debug-Bench.
Action Description Modifies State
Get_IIS Compute Irreducible Infeasible Subsystem No
Check_Slack Get constraint slack values No
Check_Bound Get variable bound status No
[…]
Drop(c)(c) Remove constraint entirely Yes
[…]
Submit Submit current model for evaluation Yes
B.2 Prompt Templates
We use three prompt variants in our experiments.
Baseline Prompt. The minimal prompt provides the problem description, code, and IIS without additional guidance:
You are an OR debugging assistant. Given an
infeasible linear program, analyze the IIS
and suggest fixes to restore feasibility.
Problem: {problem_nl}
Code: {code}
Status: INFEASIBLE
IIS: {iis_constraints}
Provide your diagnosis and suggested action.
Chain-of-Thought (CoT) Prompt. The CoT prompt adds explicit reasoning steps:
You are an OR debugging assistant. Follow
this reasoning process:
1. ANALYZE: Examine each IIS constraint’s
   role in the problem formulation
2. IDENTIFY: Determine the root cause
   constraint causing infeasibility
3. PROPOSE: Suggest a minimal fix that
   preserves problem semantics
4. VERIFY: Explain why the fix resolves
   the conflict
Problem: {problem_nl}
Code: {code}
Status: INFEASIBLE
IIS: {iis_constraints}
Optimal Workflow Prompt. Used for SFT data collection, this prompt encodes expert heuristics:
You are an expert OR debugger. Your task is:
1. Always get the IIS first
2. Identify the single most restrictive
   constraint in the IIS
3. Apply minimal relaxation (prefer relax
   over drop when possible)
4. Preserve original problem semantics
Problem: {problem_nl}
Code: {code}
Status: INFEASIBLE
IIS: {iis_constraints}
B.3 Tool Result Simulator
For training without solver access, we provide a deterministic simulator that estimates action outcomes based on ground truth labels:
Algorithm 1 Tool Result Simulator
0: Action $a$, State $s$, Ground truth $\mathcal{G}$
0: Simulated next state $s^{\prime}$
1:if $a.\text{type}\in\{\text{GET\_IIS},\text{CHECK\_SLACK}\}$ then
2:  $s^{\prime}\leftarrow s$ {Info gathering, no change}
3:else if $a.\text{target}\in\mathcal{G}.\text{key\_constraints}$ then
4:  if $a.\text{type}=\text{RELAX}$ then
5:   $s^{\prime}.\text{status}\leftarrow\text{OPTIMAL}$ w.p. 0.9
6:  else if $a.\text{type}=\text{DROP}$ then
7:   $s^{\prime}.\text{status}\leftarrow\text{OPTIMAL}$ w.p. 0.7
8:  end if
9:else if $a.\text{target}\in\mathcal{G}.\text{IIS}$ then
10:  Reduce $|\text{IIS}|$ by 1 w.p. 0.5
11:else
12:  $s^{\prime}.\text{status}\leftarrow\text{INFEASIBLE}$ {Wrong target}
13:end if
14:return $s^{\prime}$
B.4 Task Result Evaluation
We categorize episode outcomes into three classes:
  • Full Success: Status reaches Optimal and the Optimality Preservation score $\text{OP}>0.95$ (objective value within 5% of original).
  • Partial Success: Status reaches Optimal but the objective is degraded, with $0.8<\text{OP}\leq 0.95$.
  • Failure: The model remains Infeasible after max_steps, or $\text{OP}\leq 0.8$ indicating the fix changed the objective by >>20%.
The RR@k metric counts full successes achieved within $k$ steps. Partial successes are counted for the overall Recovery Rate (RR) but not for RR@k to reward efficient diagnosis.
B.5 Gurobi Configuration
We use Gurobi 11.0.0 with the following configuration:
Table 17: Gurobi solver configuration.
Parameter Value
IISMethod 1 (minimal IIS)
TimeLimit 10 seconds
OutputFlag 0 (suppress output)
Threads 4
MIPGap 0.01
Why Minimal IIS? Gurobi offers several IIS computation methods. We use the minimal method (IISMethod=1) because:
  • It produces the smallest possible IIS, making diagnosis more focused.
  • It is deterministic across runs.
  • It completes within reasonable time (typically <<1s) for our problem sizes.
B.6 Evaluation Protocol
The evaluation follows a standardized multi-attempt protocol:
Algorithm 2 Multi-Attempt Evaluation Protocol
0: Problem set $\mathcal{P}$, model $M$, max attempts $K$, max steps $T$
0: Metrics $\{\text{RR}@{1},\ldots,\text{RR}@{K},\text{DA},\text{Steps}\}$
1:for each problem $p\in\mathcal{P}$ do
2:  $\text{success}\leftarrow\text{False}$
3:  $\text{first\_success\_step}\leftarrow\infty$
4:  for $k=1$ to $K$ do
5:   Run episode with model $M$ on problem $p$
6:   if episode reaches Optimal at step $t$ then
7:    $\text{success}\leftarrow\text{True}$
8:    $\text{first\_success\_step}\leftarrow\min(\text{first\_success\_step},t)$
9:   end if
10:  end for
11:  Record success, first_success_step, DA, total_steps
12:end for
13: Compute aggregate metrics
14:return Metrics
Stratified Sampling. To ensure balanced evaluation across difficulty levels:
  • 180 samples from Types A–D (Easy)
  • 158 samples from Types E–G (Hard)
  • 112 samples from Types H–I (Expert)
Total: 450 stratified samples per model evaluation.
B.7 Reproducibility Checklist
  • $\checkmark$
    Code and data will be released upon publication
  • $\checkmark$
    Random seeds fixed for all experiments
  • $\checkmark$
    Hardware and software versions documented
  • $\checkmark$
    Training hyperparameters fully specified
  • $\checkmark$
    Evaluation protocol standardized across models
  • $\checkmark$
    Gurobi configuration deterministic
Appendix C Benchmark Construction Details
This appendix details the algorithms and validation procedures used to construct OR-Debug-Bench and OR-Bias-Bench.
C.1 Saboteur Generation
The Saboteur generates controlled infeasibilities by applying targeted corruptions to feasible LP instances. We describe two representative error types.
Type A: Direction Flip. This error reverses the sense of an inequality constraint, turning a minimum requirement into a maximum limit or vice versa:
Algorithm 3 Sabotage: Direction Flip (Type A)
0: Feasible model $M$, target constraint $c$
0: Infeasible model $M^{\prime}$, ground truth $(c,\text{``flip''})$
[…]
2:if $c.\text{sense}=\text{``}\geq\text{''}$ then
3:  $c^{\prime}.\text{sense}\leftarrow\text{``}\leq\text{''}$
4:else
5:  $c^{\prime}.\text{sense}\leftarrow\text{``}\geq\text{''}$
6:end if
7: Replace $c$ with $c^{\prime}$ in $M^{\prime}$
8:return $M^{\prime}$, $(c,\text{``flip''})$
Example. Original: $x+y\geq 10$ (minimum production requirement). Sabotaged: $x+y\leq 10$ (maximum limit, conflicts with other minimums).
Type E: Resource Over-allocation. This error increases demand constraints beyond available capacity:
Algorithm 4 Sabotage: Resource Over-allocation (Type E)
0: Feasible model $M$, demand constraints $\mathcal{D}$, factor $\alpha>1$
0: Infeasible model $M^{\prime}$, ground truth
[…]
4:assert $\sum_{c\in\mathcal{D}}c.\text{rhs}>$ capacity
5:return $M^{\prime}$, $(c^{*},\text{``reduce rhs''})$
Example. Original demands sum to 90 with capacity 100. Sabotaged demands sum to 117 ($\alpha=1.3$), exceeding capacity.
C.2 Newsvendor Generation
For OR-Bias-Bench, we generate newsvendor scenarios with controlled Critical Ratio (CR) distributions.
Algorithm 5 Newsvendor Scenario Generation
0: Target CR range $\text{CR}_{\min},\text{CR}_{\max}]$
0: Scenario parameters $(p,c,s,\mu,\sigma,Q^{*})$
[…]
8:return $(p,c,s,\mu,\sigma,Q^{*})$
The unit cost $c$ derives from the target CR using the relationship $\text{CR}=(p-c)/(p-s)$. This ensures the generated scenario has the desired CR while maintaining realistic cost structures.
C.3 Newsvendor Curriculum Design
The 4-level curriculum for OR-Bias-Bench is designed to progressively test rationality under increasing complexity.
Level Design Rationale.
  • L1 (Foundations): CR $\in[0.4,0.6]$ produces $Q^{*}\approx\mu$ (within $\pm 0.25\sigma$), minimizing pull-to-center effects. This establishes baseline formula application capability.
  • L2 (Bias Traps): CR $\in[0.05,0.2)$ yields $Q^{*}<\mu-\sigma$ (order less than mean minus one standard deviation), while CR $\in(0.8,0.95]$ yields $Q^{*}>\mu+\sigma$. These extremes maximally trigger the pull-to-center bias documented in behavioral operations research.
  • L3 (Robustness): Distractors test whether models can filter irrelevant information. Five distractor types were selected based on common supply chain context that does not affect the single-period newsvendor decision.
  • L4 (Expert): Censored demand requires inferring $\mu$ and $\sigma$ from percentiles using $\mu=P_{50}$ and $\sigma=(P_{75}-P_{25})/1.35$ for normal distributions. This tests parameter inference capability beyond formula application.
Distractor Types. Table 18 lists the five distractor categories injected in Level 3 scenarios. Each distractor provides contextually plausible but decision-irrelevant information.
Table 18: Distractor types injected in Level 3 scenarios. None affect the optimal newsvendor quantity.
Distractor Type Example Why Irrelevant
Warehouse capacity “Storage limit: 500 units” No capacity constraint in model
Competitor pricing “Competitor sells at \$45” Single-firm decision
Shelf life “Product expires in 30 days” Single-period model
Historical trends “Sales grew 10% last year” Already reflected in $\mu$, $\sigma$
Seasonal factors “Holiday season approaching” Already in demand parameters
Censored Demand (L4) Algorithm. Level 4 scenarios present demand information as percentiles rather than distribution parameters. The generation algorithm:
  1. 1.
    Generate true parameters $(\mu,\sigma)$ from standard ranges
  2. 2.
    Compute percentiles: $P_{25}=\mu+\sigma\cdot\Phi^{-1}(0.25)$, $P_{50}=\mu$, $P_{75}=\mu+\sigma\cdot\Phi^{-1}(0.75)$
  3. 3.
    Present scenario using only $(P_{25},P_{50},P_{75})$
  4. 4.
    Models must infer: $\hat{\mu}=P_{50}$, $\hat{\sigma}=(P_{75}-P_{25})/1.35$
The constant 1.35 is the interquartile range of a standard normal distribution ($\Phi^{-1}(0.75)-\Phi^{-1}(0.25)\approx 1.35$).
C.4 Stratification and ID/OOD Design
In-Distribution (ID) Set. The 400-sample ID evaluation set contains 100 samples from each curriculum level (L1–L4), ensuring balanced coverage:
  • CR distribution spans the full range $[0.05,0.95]$
  • Prompt complexity ranges from clean (L1–L2) to noisy (L3) to censored (L4)
  • Enables per-level performance analysis to diagnose specific failure modes
Out-of-Distribution (OOD) Set. The 200-sample OOD set contains only L3 and L4 scenarios (100 each), testing:
  • Robustness to distractors (L3): Can models filter irrelevant context?
  • Parameter inference capability (L4): Can models derive $(\mu,\sigma)$ from percentiles?
  • Generalization beyond clean prompts: No L1–L2 samples in OOD
Stratification Procedure. Scenarios are stratified by CR bucket to ensure no bucket is over- or under-represented:
  1. 1.
    Partition scenarios by curriculum level (L1–L4)
  2. 2.
    Within each level, bin by CR: very_low (<<0.2), low (0.2–0.4), neutral (0.4–0.6), high (0.6–0.8), very_high (>>0.8)
  3. 3.
    Sample proportionally from each bin to achieve target distribution
  4. 4.
    Verify final CR histogram matches target uniform distribution
Dataset Scale. The complete newsvendor generator produces 57,000+ scenarios across all curriculum levels and CR ranges. The released benchmark contains 2,000 instances (1,000 ID + 1,000 OOD); we evaluate on stratified subsets (400 ID + 200 OOD) ensuring:
  • Statistical power: 100+ samples per level provides reliable performance estimates
  • Diversity: All CR buckets represented to detect systematic biases
  • Discriminative power: OOD tests generalization beyond training distribution
C.5 Quality Verification
Every benchmark instance passes a four-fold validation pipeline to ensure quality.
Algorithm 6 Four-Fold Validation Pipeline
0: Original model $M$, sabotaged model $M^{\prime}$, fix $f$
0: Boolean: instance passes validation
1: {Check 1: Original feasibility}
[…]
3:if $M.\text{status}\neq\text{OPTIMAL}$ then
4:  return false
5:end if
6: {Check 2: Sabotaged infeasibility}
[…]
8:if $M^{\prime}.\text{status}\neq\text{INFEASIBLE}$ then
9:  return false
10:end if
11: {Check 3: IIS validity}
[…]
14:if $|\mathcal{I}|=0$ or $f.\text{target}\notin\mathcal{I}$ then
15:  return false
16:end if
17: {Check 4: Fix effectiveness}
[…]
20:if $M^{\prime\prime}.\text{status}\neq\text{OPTIMAL}$ then
21:  return false
22:end if
23:return true
Validation Statistics. Out of 1,200 candidate OR-Debug-Bench instances generated:
  • 98.2% passed Check 1 (original feasibility)
  • 95.7% passed Check 2 (sabotaged infeasibility)
  • 92.4% passed Check 3 (IIS contains target)
  • 89.1% passed Check 4 (fix restores optimality)
The final benchmark contains 900 instances (75% acceptance rate) that pass all four checks, ensuring every problem has a verified ground truth solution.
C.6 Error Type Examples
We provide concrete code examples for each error type, showing the original feasible formulation and the sabotaged infeasible version.
Type A: Direction Flip.
# Original (feasible)
m.addConstr(x + y >= 10, "min_production")
# Requires at least 10 units
# Sabotaged (infeasible)
m.addConstr(x + y <= 10, "min_production")
# Contradicts other minimum requirements
The direction flip creates a contradiction when combined with other constraints that require $x+y>10$.
Type B: RHS Miscalculation.
# Original (feasible)
m.addConstr(demand <= 100, "demand_limit")
# Demand cannot exceed 100
# Sabotaged (infeasible)
m.addConstr(demand <= 50, "demand_limit")
# Insufficient to meet requirements
The reduced RHS makes it impossible to satisfy downstream demand constraints.
Type C: Upper Bound Conflict.
# Original (feasible)
m.addConstr(x[0] <= 100, "capacity")
# Sabotaged (infeasible)
m.addConstr(x[0] <= 40, "capacity")
# When minimum requirement is x[0] >= 60
The tightened upper bound conflicts with a lower bound on the same variable.
Type D: Lower Bound Conflict.
# Original (feasible)
m.addConstr(x[0] >= 0, "non_negative")
# Sabotaged (infeasible)
m.addConstr(x[0] >= 100, "non_negative")
# When maximum capacity is x[0] <= 80
The inflated lower bound exceeds the variable’s upper limit.
Type E: Resource Over-allocation.
# Original (feasible)
m.addConstr(quicksum(demand[i] for i in I)
            <= 90, "total_demand")
# Sabotaged (infeasible)
m.addConstr(quicksum(demand[i] for i in I)
            <= 120, "total_demand")
# When total capacity is only 100
Individual demand constraints sum to more than available capacity.
Type F: Capacity Violation.
# Original (feasible)
m.addConstr(production <= 100, "capacity")
# Sabotaged (infeasible)
m.addConstr(production <= 60, "capacity")
# When minimum production requirement is 80
The capacity constraint is too restrictive for minimum requirements.
Type G: Flow Imbalance.
# Original (feasible)
m.addConstr(inflow[j] - outflow[j] == 0,
[…]
# Sabotaged (infeasible)
m.addConstr(inflow[j] - outflow[j] == 50,
[…]
# Creates impossible flow requirement
The modified balance constraint cannot be satisfied by any feasible flow.
Type H: Multi-constraint Conflict. Multiple constraints interact to create infeasibility:
# Sabotaged constraints
m.addConstr(x + y >= 80, "c1")
m.addConstr(x <= 30, "c2")
m.addConstr(y <= 40, "c3")
# No solution: max(x+y) = 70 < 80
Resolving the conflict requires understanding how all three constraints interact.
Type I: Composite Error. Combines multiple error types—e.g., a direction flip (Type A) that creates a cascade of bound conflicts (Type C/D):
# Multiple sabotaged constraints
m.addConstr(x + y <= 50, "flipped")    # Type A
m.addConstr(x >= 30, "bound1")          # Type D
m.addConstr(y >= 30, "bound2")          # Type D
# Infeasible: x + y >= 60 > 50
C.7 Dataset Statistics
Table 19 provides detailed statistics for both benchmarks.
Table 19: Dataset statistics for OR-Debug-Bench and OR-Bias-Bench.
Metric OR-Debug-Bench OR-Bias-Bench
Size Total pool $\scriptstyle\sim$5,000 2,000
Evaluation samples 450 600
     (ID / OOD) 400 / 200
Distribution Types A–D 44%
Types E–G 33%
Types H–I 22%
CR range [0.05, 0.95]
Complexity Avg constraints 9.9 1
Avg variables 8.4 1
Avg IIS size 4.3
C.8 Robust Injection Methods
Basic injection methods often fail when the seed problem structure does not align with the corruption strategy. We develop robust methods that adaptively select targets based on problem characteristics. This section details two representative robust methods.
Type A Robust: Slack-Based Constraint Selection. Rather than randomly selecting a constraint to flip, we solve the original problem and rank constraints by slack magnitude. Constraints with minimal slack are closest to their bounds and most likely to create infeasibility when flipped.
Algorithm 7 Robust Type A Injection: Slack-Based Selection
0: Feasible model $M$, num_candidates $k=10$
0: Infeasible model $M^{\prime}$, ground truth $(c^{*},\text{``flip''})$
1: Solve $M$, extract slack values $\{s_{c}\}$ for all constraints
2: Sort constraints by $|s_{c}|$ ascending (tightest first)
3:for $c$ in top-$k$ candidates do
4:  $M^{\prime}\leftarrow\text{flip\_direction}(M,c)$
5:  $M^{\prime}.\text{optimize}()$
6:  if $M^{\prime}.\text{status}=\textsc{Infeasible}{}$ then
7:   $M^{\prime}.\text{computeIIS}()$
8:   if $c\in\text{IIS}(M^{\prime})$ then
9:    return $M^{\prime}$, $(c,\text{``flip''})$
10:   end if
11:  end if
12:end for
13:return failure
This approach improves Type A injection success from 30% to 95%. Tightly-bound constraints have minimal slack and are most sensitive to direction changes.
Type C Robust: 4-Tier Fallback Strategy. Type C (upper bound conflict) is particularly challenging because it requires creating a conflict between an upper bound and existing constraints. We implement a cascaded fallback strategy:
Algorithm 8 Robust Type C Injection: 4-Tier Fallback
0: Feasible model $M$
0: Infeasible model $M^{\prime}$, ground truth
1: {Tier 1: High dual value targeting}
2: Solve $M$, extract dual values $\{\pi_{c}\}$
3:for $c$ in constraints with $c.\text{sense}=$$\geq$" sorted by $|\pi_{c}|$ desc do
4:  Remove positive coefficient terms from $c$
5:  if results in infeasibility with $c\in\text{IIS}$ then
6:   return success
7:  end if
8:end for
9: {Tier 2: Coefficient sign flip}
10:for $c$ in constraints with $c.\text{sense}=$$\leq$" do
11:  Flip signs of positive coefficients in $c$
12:  if results in infeasibility with $c\in\text{IIS}$ then
13:   return success
14:  end if
15:end for
16: {Tier 3: Coefficient scaling}
17:for $c$ in all inequality constraints do
18:  Scale all coefficients in $c$ by factor 10
19:  if results in infeasibility with $c\in\text{IIS}$ then
20:   return success
21:  end if
22:end for
23: {Tier 4: Guaranteed fallback}
24: Select variable $x$ with largest feasible range
25: Add tight bounds: $\leq x^{*}, xx$$\geq x^{*}+\epsilon$
26:return success (guaranteed)
The 4-tier approach increases Type C success from 0% (when Tier 1 alone fails) to 72% overall. Tier 4 provides a guaranteed fallback but produces simpler infeasibilities, so earlier tiers are preferred.
Success Rate Comparison. Table 20 compares basic and robust injection methods across all error types.
Table 20: Injection success rates: basic vs. robust methods.
Type Basic Robust Robust Method
A 30% 95% Slack-based selection
B 95% 98% RHS sensitivity analysis
C 0% 72% 4-tier fallback
D 85% 96% Bound gap targeting
E 70% 88% Capacity utilization analysis
F 65% 85% Bottleneck identification
G 60% 82% Flow balance verification
H 45% 75% Constraint interaction graph
I 40% 70% Composite strategy selection
C.9 Rejection and Regeneration Statistics
Each benchmark instance must pass a four-fold validation pipeline. Table 21 shows pass rates at each validation phase, broken down by error type.
Table 21: Validation pass rates by error type across four phases (1,200 candidate instances). Phase 1: original feasibility; Phase 2: sabotaged infeasibility; Phase 3: IIS contains target; Phase 4: fix restores optimality.
Type Phase 1 Phase 2 Phase 3 Phase 4 Final
A 100% 95% 92% 95% 82%
B 100% 100% 100% 100% 100%
C 100% 72% 68% 95% 62%
D 100% 98% 96% 98% 92%
E 100% 85% 80% 90% 72%
F 100% 82% 78% 88% 68%
G 100% 78% 75% 85% 64%
H 100% 75% 72% 82% 60%
I 100% 80% 78% 90% 70%
Overall 100% 85% 82% 91% 75%
Failure Analysis. The primary failure modes vary by error type:
  • Types A, C: Phase 2 failures occur when the flipped constraint does not interact with other constraints to create infeasibility.
  • Types E–G: Phase 3 failures occur when the IIS contains related but not the exact target constraint.
  • Types H–I: Phase 4 failures occur when the ground-truth fix does not fully restore feasibility due to cascading effects.
Regeneration Iterations. Problems failing validation are regenerated with a different seed problem or sabotage target. Table 22 shows the distribution of regeneration iterations required.
Table 22: Regeneration iterations required to pass validation.
Iterations 0 (first try) 1 2 $\geq$3
Percentage 87% 9% 3% 1%
Cumulative 87% 96% 99% 100%
87% of problems pass on first generation. Problems requiring $\geq$3 iterations (1%) are typically Type C or H, where finding a valid sabotage target is difficult.
C.10 Anti-Gaming Design Rationale
Benchmarks can inadvertently reward pattern matching over genuine reasoning. We implement several mechanisms to prevent gaming.
Why Randomized Naming Matters. In preliminary experiments with semantically-named constraints (e.g., c_key_capacity, c_target_demand), we observed that models achieved 15% higher apparent accuracy by learning to target constraints with “key” or “target” in their names. This correlation existed because our ground-truth labeling naturally assigned such names to important constraints.
By switching to UUID-based naming (e.g., c_53e476_ub, c_8a2f91_eq), we eliminate this shortcut. The naming pattern for Types G, H, and I uses:
[…]
where sense_suffix encodes only the constraint type (ub/lb/eq), not its semantic role.
Hidden Dependency Design (Type F). Type F problems are constructed so that the IIS reveals a symptom constraint $c_{s}$, but the root cause is a bound modification on variable $x$ elsewhere:
  1. 1.
    Original: $x\leq 100$ with constraint $c_{s}:\sum_{i}a_{i}x_{i}\leq b$ depending on $x$
  2. 2.
    Sabotaged: $x\leq 40$ (hidden modification) causes $c_{s}$ to become infeasible
  3. 3.
    IIS contains $c_{s}$ but not the bound on $x$
Models that blindly relax $c_{s}$ fail because the real issue is the tightened bound on $x$. Solving Type F requires reasoning about variable dependencies.
Cascading Conflict Design (Type G). Type G problems include two conflicts where the second is masked until the first is resolved:
  1. 1.
    Primary conflict: Constraint $c_{1}$ conflicts with $c_{2}$ (appears in initial IIS)
  2. 2.
    Secondary conflict: After fixing $c_{1}$, constraint $c_{3}$ conflicts with $c_{4}$
The benchmark includes 15% of problems with such structures. Models that stop after one fix fail; those that iterate diagnosis succeed.
Optimal Selection Challenge (Type I). Type I problems have multiple valid fixes, but only one preserves the optimal objective value:
  • Fix A: Relax $c_{1}$ by 10% $\rightarrow$ Feasible, objective drops 5%
  • Fix B: Relax $c_{2}$ by 5% $\rightarrow$ Feasible, objective drops 15%
  • Fix C: Modify $c_{3}$ expression $\rightarrow$ Feasible, objective preserved
The ground truth labels Fix C as correct. This tests whether models consider solution quality, not just feasibility.
C.11 Difficulty Calibration Procedure
We calibrate difficulty levels through iterative testing against a baseline SFT model. The procedure ensures each difficulty level provides meaningful differentiation.
Calibration Process.
  1. 1.
    Initial parameter selection: Set IIS size ranges based on error type complexity.
  2. 2.
    Baseline evaluation: Run SFT model on 100 problems per configuration.
  3. 3.
    Target verification: Check if RR@5 falls within target range.
  4. 4.
    Parameter adjustment: If too easy (RR@5 >> target), increase IIS size or use harder error types. If too hard (RR@5 << target), decrease IIS size or use easier error types.
  5. 5.
    Iteration: Repeat until convergence (typically 2–3 iterations).
Calibration Results. Table 23 shows the final calibrated parameters and observed SFT performance.
Table 23: Difficulty calibration results after iterative refinement.
Level IIS Range Target RR@5 Observed RR@5 Iterations
Easy 2–4 $\geq$90% 93.1% 1
Hard 5–10 70–85% 78.4% 2
Expert 8–15 45–70% 62.2% 3
Why These Ranges. The target ranges were chosen based on benchmark utility:
  • >>90% (Easy): Ensures baseline models can solve simple problems, establishing floor performance.
  • 70–85% (Hard): Provides room for improvement while remaining tractable.
  • 45–70% (Expert): Challenges advanced models without being impossibly difficult.
  • <<45%: Rejected as too difficult—problems often have ambiguous fixes or require domain knowledge beyond general OR competence.
The Expert level required 3 calibration iterations because initial IIS sizes of 10–20 resulted in only 35% RR@5, below our utility threshold. Reducing to 8–15 achieved the target range while maintaining genuine difficulty.
Appendix D MDP-Based Training Details
This appendix provides complete specifications of the MDP formulation and training procedures.
D.1 State Space
The OR-Debug-Bench environment maintains a structured state representation with eight components:
  • Problem description: Natural language specification of the optimization problem
  • Code: Current Gurobi/Pyomo model code
  • Solver status: One of OPTIMAL, INFEASIBLE, UNBOUNDED, or ERROR
  • IIS log: List of constraint names in the current Irreducible Infeasible Subsystem
  • Slack values: Constraint slack values (populated after CHECK_SLACK)
  • Bound status: Variable bound information (populated after CHECK_BOUND)
  • History: Sequence of previous actions taken in the episode
  • Step counter: Current step number (0 to max_steps)
The state is serialized to a prompt string for the LLM, including the problem description, current code, solver status, and relevant diagnostic information based on previous actions.
D.2 Action Space
Actions follow a hierarchical structure separating information gathering from model modification:
Diagnostic actions (no state change, zero cost):
  • GET_IIS: Compute the Irreducible Infeasible Subsystem
  • CHECK_SLACK: Retrieve constraint slack values
  • CHECK_BOUND: Retrieve variable bound status
Repair actions (modify model, increment step counter):
  • RELAX(constraint, delta): Increase or decrease the right-hand side by delta
  • DROP(constraint): Remove the specified constraint from the model
  • REWRITE(constraint, expr): Replace the constraint with a new expression
Meta actions:
  • SUBMIT: Submit the current model for final evaluation
  • RESTART: Reset to the initial sabotaged state
Diagnostic actions do not count toward the step limit for RR@k computation; repair and meta actions increment the step counter.
D.3 Reward Function
The composite reward function balances outcome, diagnostic accuracy, and efficiency:
Algorithm 9 Compute Reward
0: State $s$, action $a$, next state $s^{\prime}$, ground truth $\mathcal{G}$
0: Reward $r\in\mathbb{R}$
[…]
2: {Outcome Reward (50%)}
3:if $s^{\prime}.\text{status}=\text{OPTIMAL}$ then
4:  $r\leftarrow r+0.5\times 100$
5:else if $s^{\prime}.\text{status}=\text{INFEASIBLE}$ then
6:  $r\leftarrow r+0.5\times(-50)$
7:end if
8: {Diagnostic Accuracy Reward (30%)}
9:if $a$ contains diagnosis $D$ then
10:  $\text{DA}\leftarrow|D\cap\mathcal{G}.\text{IIS}|/|\mathcal{G}.\text{IIS}|$
11:  $r\leftarrow r+0.3\times(\text{DA}\times 100)$
12:end if
13: {Efficiency Reward (20%)}
[…]
16: {Faithfulness Penalty}
17:if $a.\text{type}\in\{\text{RELAX},\text{DROP},\text{REWRITE}\}$ then
18:  if $a.\text{target}\notin s^{\prime}.\text{IIS}$ then
19:   $r\leftarrow r-20$ {Penalize off-target fixes}
20:  end if
21:end if
22:return $r$
The 50%/30%/20% weighting was determined through ablation (see Appendix G). The faithfulness penalty discourages repairs that do not address the identified infeasibility source.
D.4 PRM Training Details
The Process Reward Model (PRM) provides step-level supervision for GRPO training.
Label Generation. We assign labels to each step in a trajectory based on progress indicators:
Algorithm 10 Generate Step Labels for PRM
0: Trajectory $\tau=[(s_{0},a_{0}),\ldots,(s_{T},a_{T})]$, ground truth $\mathcal{G}$
0: Labels $y_{0},\ldots,y_{T}]$
1:for $t=0$ to $T$ do
2:  if $s_{t+1}.\text{status}=\text{OPTIMAL}$ then
3:   $y_{t}\leftarrow 1.0$ {Problem solved}
4:  else if $t>0$ and $|s_{t+1}.\text{IIS}|<|s_{t}.\text{IIS}|$ then
5:   $y_{t}\leftarrow 1.0$ {IIS shrinking}
6:  else if $a_{t}.\text{diagnosis}\cap\mathcal{G}.\text{IIS}\neq\emptyset$ then
7:   $y_{t}\leftarrow 0.5$ {Correct diagnosis}
8:  else if $a_{t}.\text{type}\in\{\text{GET\_IIS},\text{CHECK\_SLACK}\}$ then
9:   $y_{t}\leftarrow 0.2$ {Information gathering}
10:  else
11:   $y_{t}\leftarrow 0.0$ {No progress}
12:  end if
13:end for
14:return $y_{0},\ldots,y_{T}]$
Training Configuration.
Table 24: PRM training hyperparameters.
Parameter Value
Base model Qwen/Qwen3-8B
Method LoRA ($r=8$, $\alpha=16$)
Epochs 3
Batch size 8
Learning rate $2\times 10^{-5}$
Warmup ratio 0.1
Metric for best model AUC-ROC
The PRM achieves AUC-ROC of 0.94 on held-out step labels (309 test samples from 1,548 total labels), with correlation metrics: Pearson r=0.87 (p¡1e-90), Spearman r=0.82 (p¡1e-70). This indicates good discrimination between productive (avg score 0.72) and unproductive steps (avg score 0.50).
D.5 GRPO Training Curves
Table 25 shows the evolution of key metrics during GRPO training.
Table 25: GRPO training metrics by epoch.
Epoch Mean Reward Std Reward RR@5 (val) DA (val)
1 45.2 28.3 92.0% 56.1%
2 62.8 22.1 93.5% 58.4%
3 71.4 18.7 94.5% 60.8%
4 74.2 16.9 95.0% 62.1%
Convergence Analysis.
  • Reward variance decreases from 28.3 to 16.9, indicating policy stabilization as the model converges to consistent behavior.
  • RR@5 plateaus after epoch 4, with diminishing returns beyond this point. We select the epoch-4 checkpoint for evaluation.
  • DA improves more slowly than RR@5, confirming that learning accurate diagnosis is harder than achieving feasibility. This motivates the 30% weight on diagnostic accuracy in the reward function.
D.6 Curriculum Training Configuration
For OR-Bias-Bench, we use a three-stage curriculum over the Critical Ratio (CR) distribution:
Table 26: Curriculum stages for OR-Bias-Bench training.
Stage CR Range Samples Purpose
1 $\{0.1,0.9\}$ 200 Learn extreme directions
2 $0.15,0.25]\cup[0.75,0.85]$ 300 Calibrate boundaries
3 $[0.2,0.8]$ 400 Full distribution coverage
Stage 1 trains on extreme CR values where the optimal direction is unambiguous (high CR $\rightarrow$ order more, low CR $\rightarrow$ order less). Stage 2 refines decision boundaries. Stage 3 ensures generalization across the full distribution.
[…]
D.7 Complete Hyperparameter Tables
Table 27 and Table 28 provide complete training configurations.
Table 27: SFT training hyperparameters.
Parameter Value
Base model Qwen/Qwen3-8B
Method LoRA ($r=16$, $\alpha=32$)
Epochs 3
Batch size 4 (per GPU)
Gradient accumulation 4
Learning rate $2\times 10^{-5}$
LR scheduler Cosine
Warmup ratio 0.03
Max sequence length 4096
Weight decay 0.01
Optimizer AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.999$)
Table 28: GRPO training hyperparameters.
Parameter Value
Base model Qwen3-8B-SFT (from SFT)
Method LoRA ($r=16$, $\alpha=32$)
Epochs 4
Group size 4 (samples per prompt)
Learning rate $5\times 10^{-6}$
KL coefficient $\beta$ 0 (removed)
Clip range $\epsilon$ [0.2, 0.28] (asymmetric)
Reward components 50%/30%/20%
Max steps per episode 50
GPU memory 2 $\times$ A100 80GB
D.8 Hardware and Software Environment
All experiments were conducted on the following setup:
Table 29: Hardware and software configuration.
Component Specification
GPU 2 $\times$ NVIDIA A100 80GB
CPU AMD EPYC 7V13 (64 cores)
Memory 256 GB
Storage NVMe SSD
CUDA 12.9.1
PyTorch 2.9.1
Transformers 4.44.0
TRL 0.24.0
SGLang 0.3.5
Gurobi 11.0.0
D.9 Inference Configuration
For evaluation, we use SGLang for efficient batched inference:
Table 30: Inference configuration.
Parameter Value
Tensor parallel size 2
Batch size 16 (concurrent requests)
Max new tokens 2048
Temperature 0.0 (deterministic)
Attention backend Triton
Sampling backend PyTorch
D.10 Training Data Examples
We provide additional examples of the training data format.
SFT Data Format. Each SFT example contains an instruction, input state, and target output:
  "instruction": "Debug the infeasible
                  optimization model.",
  "input": "Problem: mip_typeA_008
    Status: INFEASIBLE
    IIS: [’c0’, ’c1’, ’c_key_upper’]",
  "output": "<think>
    IIS contains 3 constraints.
    c_key_upper limits x <= 50 but
    c0 and c1 require x >= 80.
    </think>
    Action: RELAX(c_key_upper, 30)"
GRPO Prompt Format. GRPO training uses prompts without target outputs:
  "prompt": "Debug the infeasible model.
    Problem: mip_typeE_042
    Status: INFEASIBLE
    IIS: [’supply’, ’d1’, ’d2’, ’d3’]",
  "ground_truth": {
    "key_constraint": "d2",
    "expected_fix": "RELAX(d2, 15)"
The model generates completions, which are scored using the composite reward function.
OR-Bias-Bench Training Format. Newsvendor scenarios include all parameters:
  "scenario": {
    "price": 50, "cost": 35,
    "salvage": 10, "mean": 100,
    "std": 20, "CR": 0.375
  "optimal_Q": 93.6,
  "stage": 1  # Curriculum stage
Appendix E OR-Debug-Bench: Complete Results
E.1 Main Results
Table 31 shows complete OR-Debug-Bench results for all 26 models evaluated. All models report RR@5 and DA from multi-attempt evaluation.
Table 31: Complete OR-Debug-Bench results (450 stratified samples per model). All models evaluated with multi-attempt protocol.
Model Type $R$ RR@5 DA Steps
Qwen3-8B-GRPO Local 100% 95.3% 62.4% 2.25
Qwen3-8B-Curriculum Local 100% 94.0% 61.7% 2.22
Qwen3-8B-DAPO Local 100% 93.8% 60.4% 2.31
Qwen3-8B-SFT Local 99.8% 93.1% 60.8% 2.34
claude-sonnet-4 API 100% 86.2% 50.1% 3.71
claude-haiku-4.5 API 99.3% 86.0% 53.1% 3.89
o4-mini API 97.8% 86.2% 47.8% 3.15
o1 API 99.8% 82.9% 47.8% 3.78
gpt-5.2-chat API 99.8% 81.8% 40.9% 3.72
qwen2.5-7b API 97.8% 77.8% 40.1% 4.36
claude-opus-4 API 94.2% 76.9% 49.0% 3.92
o3 API 96.7% 75.8% 50.9% 4.23
DeepSeek-V3.2 API 99.3% 58.9% 44.8% 4.86
gpt-4.1 API 94.4% 71.6% 36.2% 4.41
gemini-2.5-flash API 84.2% 70.7% 19.2% 3.23
Llama-3.3-70B API 93.8% 60.9% 46.9% 4.81
gpt-5-mini API 100% 66.9% 37.6% 4.74
gemini-2.5-pro API 62.9% 62.7% 52.5% 0.83
qwen2.5-32b API 98.9% 61.1% 32.0% 4.98
DeepSeek-R1 API 99.1% 56.7% 34.5% 5.08
qwen2.5-max API 99.1% 54.9% 42.6% 5.97
qwen2.5-14b API 88.9% 53.6% 35.4% 6.32
gemini-2.0-flash API 85.6% 52.4% 18.3% 5.63
gpt-4.1-mini API 93.1% 49.8% 26.0% 6.04
kimi-k2 API 55.3% 40.4% 25.6% 2.48
claude-3.7-sonnet API 98.9% 31.6% 46.8% 8.06
E.2 Per-Error-Type Analysis
Per-error-type analysis is available upon request. Table 31 provides results for all 26 models.
E.3 Token Efficiency
Table 32 shows representative token usage for selected models.
Table 32: Token usage per episode on OR-Debug-Bench (representative models).
Model Tokens Steps Tok/Success
Qwen3-8B-GRPO 2,011 2.25 2,109
Qwen3-8B-SFT 2,103 2.34 2,259
claude-sonnet-4 5,417 3.71 6,283
o4-mini 6,246 3.36 7,985
gpt-4.1 7,260 4.78 11,503
gpt-5-mini 10,544 5.12 18,110
Efficiency Observations.
[…]
  • Step count correlates with token usage ($r=0.79$), but response length per step varies by model.
E.4 RAG Ablation
Table 33 presents the complete RAG ablation across retrieval strategies and $k$ values.
Table 33: RAG ablation on OR-Debug-Bench (200 samples).
Configuration $R$ RR@5 DA Steps
No RAG (baseline) 99.8% 83.0% 80.0% 3.26
quick_fix (kk=3) 99.5% 80.0% 66.6% 2.58
reasoning (kk=3) 100% 93.5% 51.8% 1.60
by_type (kk=1) 99.5% 86.5% 80.0% 2.19
by_type (kk=3) 100% 94.5% 82.0% 1.51
by_type (kk=5) 100% 96.5% 85.0% 1.62
by_type (kk=7) 100% 97.0% 85.0% 1.53
Retrieval Strategy Comparison.
  • by_type: Best overall. Retrieves cases with similar error types, providing relevant examples without giving away the solution.
  • reasoning: High RR@5 but lower DA. Provides complete reasoning chains that models copy, achieving correct fixes without learning to diagnose.
  • quick_fix: Worst performance. Too shallow for complex errors, often omitting required diagnostic steps.
$k$ Value Analysis. Performance improves from kk=1 to kk=5, then plateaus. We recommend kk=5 as the default, balancing accuracy (+13.5% over baseline) with retrieval cost.
E.5 Failure Analysis
Common Error Patterns in Failed Episodes. We analyzed 100 randomly sampled failures from Qwen3-8B-GRPO:
Table 34: Failure pattern distribution for Qwen3-8B-GRPO.
Pattern Count Example
Wrong constraint identified 22 Relaxed c3 instead of c5
Insufficient relaxation 18 Relaxed by 5, needed 10
Cascading failure 42 Fix c1 $\rightarrow$ new IIS with c7
Timeout on complex IIS 11 12+ constraints in IIS
Objective degradation 7 Fix valid but $\text{OP}<0.8$
Cascade Failure Analysis. The most common failure mode (42%) involves cascading errors where fixing one constraint reveals another. These occur predominantly on Type H–I problems:
Step 1: IIS = {c1, c3, c5}
        Action: RELAX(c1, 10)
Step 2: IIS = {c3, c7, c9}  # New conflict!
        Action: RELAX(c7, 5)
Step 3: IIS = {c5, c9, c11}  # Another new conflict
        ...continues until timeout
This pattern suggests the need for lookahead reasoning about constraint dependencies.
Appendix F OR-Bias-Bench: Complete Results
Table 35 provides complete OR-Bias-Bench results for all 24 evaluated models, reporting both rationality (valid numerical responses) and bias (deviation from rational ordering) metrics across ID and OOD splits.
Table 35: Complete OR-Bias-Bench results with ID/OOD breakdown (24 models, sorted by ID Bias).
Rationality Bias
Model ID OOD ID OOD $\Delta$
claude-haiku-4.5 99.9% 99.9% 0.0% 3.6% +3.6%
o3 93.1% 97.7% 0.4% 24.5% +24.1%
qwen2.5-max 99.4% 98.5% 0.5% 25.0% +24.5%
gpt-5-mini 99.6% 99.7% 1.2% 53.3% +52.1%
claude-sonnet-4 89.0% 93.5% 1.5% 7.7% +6.2%
gpt-4.1-mini 99.6% 99.9% 4.1% 12.1% +8.0%
Qwen3-8B-OM-SFT 99.8% 99.6% 4.9% 11.5% +6.6%
gemini-2.0-flash 97.6% 98.7% 5.9% 0.0% -5.9%
qwen2.5-14b 99.1% 99.8% 5.9% 10.7% +4.8%
o4-mini 98.5% 99.4% 6.7% 7.7% +1.0%
qwen2.5-7b 98.1% 99.3% 7.4% 8.2% +0.8%
kimi-k2 92.8% 97.6% 8.9% 2.3% -6.6%
claude-3.7-sonnet 99.8% 99.4% 11.3% 18.1% +6.8%
gpt-4.1 99.9% 100.0% 11.5% 14.6% +3.0%
qwen2.5-32b 95.9% 99.8% 15.4% 5.0% -10.4%
DeepSeek-V3.2 100.0% 99.9% 18.2% 11.3% -6.9%
gemini-2.5-flash 96.2% 98.1% 18.2% 71.2% +53.0%
claude-opus-4 92.5% 94.8% 19.0% 15.9% -3.1%
Llama-3.3-70B 92.0% 96.9% 19.2% 12.5% -6.7%
Qwen3-8B-OM-Curriculum 99.9% 99.6% 20.0% 10.4% -9.6%
o1 98.2% 98.6% 23.2% 43.6% +20.3%
DeepSeek-R1 98.6% 99.6% 44.1% 40.9% -3.2%
Qwen3-8B-OM-GRPO 96.1% 98.4% 48.0% 33.8% -14.2%
gemini-2.5-pro 98.2% 99.6% 97.9% 100.0% +2.1%
F.1 Rationality Analysis
Most models achieve >>99% rationality, consistently producing valid numerical orderings. Two exceptions stand out: o3 (93.1% ID) and claude-sonnet-4 (89.0% ID). These reasoning-heavy models occasionally produce malformed outputs when over-analyzing simple ranking tasks.
F.2 ID Bias Patterns
ID bias spans from 0.0% (claude-haiku-4.5) to 97.9% (gemini-2.5-pro):
  • Near-zero bias (<<2%): claude-haiku-4.5, o3, qwen2.5-max, gpt-5-mini—these models correctly apply EOQ/newsvendor logic on ID distributions.
  • Moderate bias (5–20%): Most API models fall here, having partially but not fully internalized OR principles.
  • High bias (>>40%): DeepSeek-R1 (44.1%), Qwen3-8B-OM-GRPO (48.0%), gemini-2.5-pro (97.9%)—these models systematically deviate from rational orderings.
F.3 OOD Generalization
[…]
Catastrophic Degradation. gpt-5-mini shows the most severe drift (+52.1%, from 1.2% to 53.3%): low ID bias does not guarantee OOD generalization. gemini-2.5-flash degrades similarly (+53.0%). These models likely memorize ID patterns rather than learn underlying principles.
Stable Performance. o4-mini (+1.0%), qwen2.5-7b (+0.8%), and gpt-4.1 (+3.0%) maintain consistent bias across distributions. These models appear to have internalized OR principles more robustly.
OOD Improvement. Curriculum training achieves the only substantial OOD improvement among trained models (-9.6%, from 20.0% to 10.4%). Other improving models include kimi-k2 (-6.6%), gemini-2.0-flash (-5.9%), and qwen2.5-32b (-10.4%).
F.4 Local Model Comparison
The three Qwen3-8B-OM variants show distinct trade-offs:
  • SFT: Best ID bias (4.9%) among local models, moderate OOD drift (+6.6%).
  • Curriculum: Higher ID bias (20.0%) but best OOD performance (10.4%, -9.6% improvement).
  • GRPO: Highest bias on both splits (48.0% ID, 33.8% OOD). Outcome-focused RL does not transfer well to bias mitigation.
This pattern confirms that curriculum training prioritizes generalization over ID memorization—critical for real-world OR applications where distribution shift is common.
Appendix G Training Ablation Studies
G.1 Reward Weight Ablation (OR-Debug)
Table 36 shows the effect of varying the diagnostic reward weight in the composite reward function.
Table 36: Reward weight ablation on OR-Debug-Bench validation set.
Diagnostic Weight RR@5 DA Steps
20% 94.8% 54.2% 2.18
30% 95.1% 58.6% 2.21
40% 95.3% 62.4% 2.25
50% 94.6% 64.1% 2.42
60% 93.2% 65.3% 2.78
Selected for final training based on the RR@5/DA trade-off. The 30–40% range achieves optimal balance; we use 30% in our final 50%/30%/20% (outcome/diagnosis/efficiency) configuration.
Reducing diagnostic weight below 30% leads to repairs that achieve feasibility without correctly identifying the root cause (low DA). Weights above 50% slow convergence by over-penalizing exploratory actions, as shown by increased step counts.
G.2 Curriculum Stage Ablation (OR-Bias)
Table 37 compares different curriculum configurations for OR-Bias-Bench.
Table 37: Curriculum ablation on OR-Bias-Bench OOD set.
[…]
No curriculum (SFT only) 11.5% +6.6%
Stage 1 only (extreme CR) 15.2% +2.1%
Stages 1+2 12.1% -3.4%
Full curriculum (1+2+3) 10.4% -9.6%
The full three-stage curriculum achieves the best OOD generalization, with each stage contributing to the final performance.
Appendix H Token Efficiency and Difficulty Analysis
This appendix presents analysis of token efficiency and problem difficulty patterns across models.
H.1 Token Efficiency
Token efficiency analysis is summarized in Table 32 (Appendix E.3).
Key Observations.
  • Token efficiency: Local models require 2,000–2,100 tokens per episode compared to 2,900–17,300 for API models. Tokens-per-success ranges from 2,109 (Qwen3-8B-GRPO) to 54,847 (claude-3.7-sonnet).
  • Step efficiency: Local models solve problems in 2.2–2.3 steps on average, while API models range from 0.7 (DeepSeek-R1) to 8.1 steps (claude-3.7-sonnet).
  • Correlation: Step count correlates with token usage ($r=0.79$), though response length per step varies widely across models (500–3,000 tokens).
H.2 Test-Time Compute Analysis
Token efficiency data is summarized in Table 32 (Appendix E.3). Key findings:
Efficiency Metrics.
[…]
H.3 Scaling with Problem Difficulty
Based on per-error-type results from the main evaluation, we group error types by empirical difficulty:
Table 38: Difficulty grouping based on type-averaged RR@5 across all 26 models.
Difficulty Error Types Avg RR@5 Characteristics
Easy B (87%), C (76%) 81.4% Clear diagnostic signals
Medium H (65%), I (64%), E (61%), D (57%) 61.6% Multiple constraint interactions
Hard A (54%), F (46%), G (45%) 48.1% Semantic reasoning required
Observations.
  • Type B is easiest, not Type A: Bound violations (Type B) produce clear diagnostic signals with 87% average success. Type A (objective direction), despite being a single-constraint error, averages only 54%—models struggle with objective function semantics.
  • Flow and capacity are hardest: Types F (flow balance, 46%) and G (capacity violations, 45%) require reasoning about multiple interacting constraints, explaining the difficulty gap.
  • Local models excel uniformly: Qwen3-8B variants achieve >>86% on all types including the hardest (F: 94–100%, G: 86–98%), while API models show type-specific weaknesses.
H.4 Recommendations for Practitioners
Based on our analysis, we provide the following recommendations:
  1. 1.
    Apply difficulty-adaptive resources: Allocate more compute to Types A, F, G (Hard, 48% avg RR@5) than Types B, C (Easy, 81% avg).
  2. 2.
[…]
  3. 3.
[…]
Appendix I Base Model Selection Study
This appendix documents our pilot study for selecting the foundation model for domain-specific training, and provides analysis of why standard prompting approaches fail on the OR debugging task.
I.1 Candidate Model Screening
We evaluated Qwen3-8B-Instruct as the foundation model for domain-specific training. Selection criteria included: (1) base performance on OR debugging, (2) improvement potential with SFT, and (3) inference efficiency.
Table 39: Foundation model screening on OR-Debug-Bench validation set (100 samples).
Model Params Base RR@5 +SFT RR@5 $\Delta$ Tokens/ep
Qwen3-8B-Instruct 8B 51.2% 93.1% +41.9% 2,100
Key Findings.
  • Base performance: Qwen3-8B achieves 51.2% RR@5 without any domain-specific training, demonstrating reasonable out-of-the-box capability for structured reasoning.
  • SFT improvement: Qwen3-8B improves by +41.9% with SFT, indicating high receptivity to domain adaptation.
  • Efficiency: Qwen3-8B generates 2,100 tokens per episode, providing efficient inference for iterative debugging.
Selection Rationale. We selected Qwen3-8B-Instruct as the foundation model based on three factors:
  1. 1.
    Strong post-SFT performance: 93.1% RR@5 after SFT demonstrates successful domain adaptation.
  2. 2.
    Good improvement potential: +41.9% delta suggests the model effectively learns from demonstration data.
  3. 3.
    Practical efficiency: Reasonable token footprint reduces training and inference costs.
I.2 Negative Result: Standard Prompting Approaches Fail
Before investing in domain-specific training, we evaluated whether standard prompting approaches could achieve competitive performance on OR-Debug-Bench. This section documents these negative results.
I.2.1 Zero-Shot Chain-of-Thought
We evaluated zero-shot CoT prompting with the instruction: “Let’s think step by step about how to debug this infeasible model.”
Table 40: Zero-shot CoT performance on OR-Debug-Bench (200 samples).
Model RR@5 DA Avg Steps Notes
gpt-5.2-chat + CoT 38.5% 22.1% 6.8 Verbose, unfocused
o4-mini + CoT 41.2% 28.4% 5.9 Better structure
Qwen3-8B + CoT 23.0% 15.6% 7.2 Often loops
Why Zero-Shot CoT Fails.
  • No feedback loop: CoT generates a single reasoning chain without iterating based on solver output. Models attempt repairs without verifying whether they resolved the infeasibility.
  • Generic reasoning patterns: CoT prompting elicits general problem-solving steps (“identify the issue, propose a solution, verify”) that lack domain-specific diagnostic actions like GET_IIS.
  • Premature commitment: Models commit to repair strategies early in the chain without exploring the constraint structure, leading to suboptimal fixes.
I.2.2 Few-Shot In-Context Learning
We evaluated 1-shot and 3-shot ICL with curated examples of successful debugging trajectories.
Table 41: Few-shot ICL performance on OR-Debug-Bench (200 samples).
Configuration RR@5 DA Avg Steps
gpt-5.2-chat (0-shot) 38.5% 22.1% 6.8
gpt-5.2-chat (1-shot) 52.3% 35.2% 4.9
gpt-5.2-chat (3-shot) 58.1% 41.6% 4.2
Qwen3-8B (0-shot) 23.0% 15.6% 7.2
Qwen3-8B (1-shot) 31.4% 24.3% 6.1
Qwen3-8B (3-shot) 38.7% 29.8% 5.4
Why Few-Shot ICL Is Insufficient.
  • Limited generalization: 3-shot ICL improves performance by +19.6% for gpt-5.2-chat but still falls far short of SFT (+41.9% for Qwen3-8B).
  • Context length constraints: Each debugging trajectory requires 500–1000 tokens. With 3 examples, the prompt consumes 1.5–3K tokens, limiting remaining context for the actual problem.
  • Example selection sensitivity: Performance depends on example choice. Random examples achieve only 48.2% RR@5, while carefully selected examples reach 58.1%—but this selection requires domain expertise.
I.2.3 Comparison Summary
Table 42 summarizes why domain-specific training outperforms prompting approaches by 54+ percentage points.
Table 42: Comparison of approaches on OR-Debug-Bench (Qwen3-8B base).
Approach RR@5 DA Gap to SFT
Zero-shot 18.4% 12.3% -74.7%
Zero-shot + CoT 23.0% 15.6% -70.1%
1-shot ICL 31.4% 24.3% -61.7%
3-shot ICL 38.7% 29.8% -54.4%
SFT 93.1% 60.8%
SFT + GRPO 95.3% 62.4% +2.2%
Implication. The 54-point gap between 3-shot ICL (38.7%) and SFT (93.1%) demonstrates that OR debugging cannot be solved through prompting alone. The task requires:
  1. 1.
    Iterative interaction: Learning to use solver feedback across multiple turns.
  2. 2.
    Domain-specific actions: Acquiring the diagnostic vocabulary (GET_IIS, CHECK_SLACK).
  3. 3.
    Strategy patterns: Learning when to diagnose vs when to repair, and how to calibrate repair magnitudes.
These capabilities cannot be induced through few-shot examples alone.
I.3 Analysis: What Makes OR Debugging Hard for Prompting?
We identify three structural properties of OR debugging that make it resistant to prompting-based solutions:
1. Multi-Turn Dependency. Unlike single-turn tasks where CoT can decompose reasoning, OR debugging requires acting on solver feedback across multiple turns. The optimal action at step $t$ depends on the solver response at step t1t-1, which cannot be simulated within a single prompt.
2. Precise Action Syntax. The action space requires exact syntax (e.g., RELAX(c_key_upper, 30)). Small errors in constraint names or numeric values lead to failed repairs. This precision requirement exceeds what few-shot examples can reliably demonstrate.
3. State-Dependent Strategy. The optimal strategy varies with problem structure:
  • Small IIS (2–3 constraints): Direct repair often succeeds.
  • Medium IIS (4–7 constraints): Diagnosis before repair improves success rate.
  • Large IIS (8+ constraints): Systematic decomposition is required.
Few-shot prompting cannot convey these conditional strategies without extensive examples that exceed context limits.
I.4 Implications for Future Work
Our negative results suggest several directions for improving prompting-based approaches:
  1. 1.
    Tool-augmented prompting: Providing models with explicit solver interfaces (rather than expecting them to generate action syntax) may reduce syntax errors.
  2. 2.
    Retrieval-augmented generation: Our RAG experiments (Appendix E.4) show that retrieving similar solved cases improves RR@5 by +13.5%, partially closing the gap to SFT.
  3. 3.
    Multi-turn demonstration: Future work could explore demonstration formats that explicitly show the feedback loop across turns, though this faces context length challenges.
These approaches address symptoms rather than the fundamental issue: prompting cannot instill the procedural knowledge that SFT provides through gradient-based learning on hundreds of examples.